# Design of Bayesian A/B Tests Controlling False Discovery Rates and Power

### v4

@@ 3 A Proxy to the Joint Sampling Distribution
4 Economical Assessment of the FDR and Average Power
We generalize the results from Theorem 1 to develop the procedure in Algorithm 2. This procedure allows practitioners to efficiently explore the ($n,\boldsymbol{\gamma}$)-space to find the $(n,\boldsymbol{\gamma})$ combination under a given optimization scheme that minimizes the sample size while satisfying criteria for the FDR and average power. Algorithm 2 is economical because we estimate the true joint sampling distribution of posterior probabilities via simulation at only two sample sizes: $n_{0}$ and $n_{1}$. The initial sample size $n_{0}$ is an input for Algorithm 2, and it can be selected based on the anticipated budget or timeline for the A/B test. We note that Algorithm 2 details a general application of our methodology, and we later describe potential modifications.
@@ 4 Economical Assessment of the FDR and Average Power
1:procedure Optimize($f(\cdot)$, $g(\cdot)$, $\boldsymbol{\delta}_{L}$, $\boldsymbol{\delta}_{U}$, $p(\boldsymbol{\eta})$, $c$, $q$, $\beta$, $m$, $\Psi$, $n_{0}$)
@@
6:     Find $n_{1}$, the smallest $n$ such that Algorithm 1 with the expits of $\{\{\hat{l}^{{}_{(n)}}_{r,k}\}_{k=1}^{K}\}_{r=1}^{m}$ yields average power      $\geq 1-\beta$, where $\boldsymbol{\gamma}_{n}$ are the optimal thresholds such that $\widehat{\text{FDR}}$ $\leq q$
7:     Compute $\{\{\widehat{Pr}(H_{1,k}~{}|~{}\boldsymbol{y}^{{}_{(n_{1})}}_{r})\}_{k=1}^{K}\}_{r=1}^{m}$ via Algorithm 1 and their logits $\{\{\hat{l}^{{}_{(n_{1})}}_{r,k}\}_{k=1}^{K}\}_{r=1}^{m}$
8:     for $d$ in $1$:$m$ do
9:         for $k$ in $1$:$K$ do
10:              Let $\hat{l}^{{}_{(n_{0})}}_{d,k}$ and $\hat{l}^{{}_{(n_{1})}}_{d,k}$ be the $d^{\text{th}}$ order statistics of $\{\hat{l}^{{}_{(n_{0})}}_{r,k}\}_{r=1}^{m}$ and $\{\hat{l}^{{}_{(n_{1})}}_{r,k}\}_{r=1}^{m}$
11:              Let $r$ be the index of the sample $\boldsymbol{y}^{{}_{(n_{1})}}_{r}$ corresponding to $\hat{l}^{{}_{(n_{1})}}_{d,k}$
12:              Now use the line $\hat{L}^{{}_{(n)}}_{r,k}$ passing through $(n_{0},\hat{l}^{{}_{(n_{0})}}_{d,k})$ and $(n_{1},\hat{l}^{{}_{(n_{1})}}_{d,k})$ to get $\hat{l}^{{}_{(n)}}_{r,k}$ for other $n$
13:     Find $n_{2}$, the smallest $n$ such that Algorithm 1 with the expits of $\{\{\hat{l}^{{}_{(n)}}_{r,k}\}_{k=1}^{K}\}_{r=1}^{m}$ yields average power      $\geq 1-\beta$, where $\boldsymbol{\gamma}_{n}$ are the optimal thresholds such that $\widehat{\text{FDR}}$ $\leq q$
14:     return $n_{2}$ as recommended $n$ and $\boldsymbol{\gamma}_{n_{2}}$ as $\boldsymbol{\gamma}$
@@
We choose the next sample size $n_{1}$ at which to estimate posterior probabilities in Line 6 of Algorithm 2. Recall that the optimal decision thresholds $\boldsymbol{\gamma}$ are linked with $n$ as described in Section 2. To account for this phenomenon, we run an optimization scheme to find the optimal $\boldsymbol{\gamma}$ for which $\widehat{\text{FDR}}$, the estimate of the FDR defined in (1), is at most $q$. This optimization scheme can be as simple or as complex as the practitioner desires; we elaborate on several potential schemes in Section 5. Line 6 introduces a subscript $n$ for the decision thresholds to emphasize this dependence on the sample size. Given $\boldsymbol{\gamma}_{n}$ for a particular $n$, we can determine whether this sample size achieves an average power of at least $1-\beta$. The sample size $n_{1}$ is the smallest $n$ such that the average power is sufficiently large. We recommend finding this sample size using binary search. We emphasize that a modified version of Algorithm 1 based on the linear approximations from Line 5 of Algorithm 2 can be efficiently implemented for new sample sizes without computing additional posterior probabilities.
@@
In Lines 10 to 12 of Algorithm 2, we construct linear approximations to logits of posterior probabilities that are less reliant on large-sample results. These approximations are obtained separately for each metric $k$ by fitting a line through the same order statistic of the marginal sampling distribution at the sample sizes $n_{0}$ and $n_{1}$. Because Theorem 1 ensures that the limiting slopes of $\text{logit}\left(p^{{}_{(n)}}_{r,k}\right)$ and $\text{logit}\left(p^{{}_{(n)}}_{r,k}~{}|~{}\{p^{{}_{(n)}}_{r,h}\}_{h=1}^{k-1}\right)$ from the proxy sampling distribution are the same, we use the marginal sampling distributions to estimate slopes for the conditional logits. To maintain the proper level of dependence in the joint sampling distribution of posterior probabilities, we group the lines $\hat{L}^{{}_{(n)}}_{r,k}$ from Line 12 across metrics based on the sample $\boldsymbol{y}^{{}_{(n_{1})}}_{r}$ that defines the linear approximations. Given the linear trend in the proxy sampling distribution quantiles discussed in Section 3, it is reasonable to construct these linear approximations based on order statistics of estimates of the true sampling distributions when the $\boldsymbol{\theta}_{r}$ value is similar for all $\boldsymbol{\eta}_{r}\sim\Psi$.
@@
However, we recommend defining $\Psi$ as a mixture of various submodels, each of which gives rise to a different $\boldsymbol{\theta}_{r}$ value, to accommodate uncertainty about which hypotheses are false. This submodel approach is demonstrated in Section 5. In that case, the process in Lines 10 to 12 can be modified. We instead split the logits of the posterior probabilities for each $n$ value into subgroups based on their submodels in $\Psi$ before constructing the linear approximations. We use this set of linear approximations to obtain the optimal sample size $n_{2}$ and its associated decision thresholds in Line 13. As a result of Theorem 1 and the total variation distance between the proxy and true sampling distributions converging to 0 as $n\rightarrow\infty$, Algorithm 2 guarantees the FDR control in (2). In Section 5, we consider the performance of this algorithm for an example based on a real A/B test involving 5 metrics.
@@ 5 Optimizely Example
We choose each set of multinomial probabilities for both groups using linear programming (Guenin et al., 2014). Linear programming allows us to construct multinomial models that satisfy the following objectives. First, the probabilities $\sum_{v=1}^{13}\eta_{j,v}$ must sum to 1 in both groups. Second, we must attain specified values for the marginal probabilities $\{\pi_{A,k}\}_{k=1}^{5}$ in group A. Third, we need the marginal probabilities $\{\pi_{B,k}\}_{k=1}^{5}$ in group B to achieve targets for lift based on anticipated values for $\{\theta_{k}\}_{k=1}^{5}$. These anticipated values determine which hypotheses in $\{H_{1,k}\}_{k=1}^{5}$ are true along with the effect sizes corresponding to the true hypotheses. We elaborate on this method to select a set of multinomial models in Appendix C.1 of the supplement.
@@
For a general A/B test, there are ($K\choose k$ ways to classify $k$ of the $K$ hypotheses as false. In this Optimizely experiment, there are $\sum_{k=1}^{4}{5\choose k}=30$ unique classifications of $\{H_{1,k}\}_{k=1}^{5}$ such that some but not all of the hypotheses are false. We choose to define $\Psi$ by specifying 30 sets of multinomial models in Appendix C.1. We index these submodels by introducing a subscript to the notation $\Psi_{\boldsymbol{\Lambda}}$, where $\varnothing\subsetneq\boldsymbol{\Lambda}\subsetneq\{1,2,3,4,5\}$ is the set of hypotheses that are false. For the submodel $\Psi_{\boldsymbol{\Lambda}}$, the multinomial models for each group ensure that $\theta_{k}=0$ for $k\in\boldsymbol{\Lambda}$ and $\theta_{k}=10\%$ otherwise. It follows that $H_{1,k}$ is false in this submodel if and only if $k\in\boldsymbol{\Lambda}$. For illustration, we define the model $\Psi$ input into Algorithm 2 as an equally weighted mixture of these 30 submodels.
We now apply the method in Algorithm 2 with balanced sample sizes ($c=1$) along with criteria for the FDR and average power defined by $q=0.05$ and $\beta=0.2$. For now, we find optimal decision thresholds under the constraint that $\boldsymbol{\gamma}=\gamma\times\mathbf{1}_{K}$. This constraint gives rise to a simple optimization scheme to select $\gamma$: the optimal decision threshold is the smallest probability such that $\widehat{\text{FDR}}\leq q$. We choose an initial sample size of $n_{0}=1.2\times 10^{4}$ using information from Optimizely’s 2012 experiment (Siroker and Koomen, 2013). We implemented our method with $10^{3}$ simulation repetitions for each of the 30 submodels, giving rise to a total of $m=3\times 10^{4}$ simulation repetitions. The prior $p(\boldsymbol{\eta})$ is comprised of independent, diffuse Dirichlet$(\mathbf{1}_{13})$ priors for the multinomial probabilities $\boldsymbol{\eta}_{j}$ in each group. These distributions are conjugate priors for $\boldsymbol{\eta}_{A}$ and $\boldsymbol{\eta}_{B}$, which we use to estimate the joint posterior of $\boldsymbol{\theta}$ via simulation.
@@
Algorithm 2 with these inputs returned a recommended $(n_{A},\gamma)$ combination of (13328,(13328,( 13328 , 0.9411)0.9411)0.9411 ). All sample size recommendations in this section pertain to the sample size per group (i.e., $n_{A}=n_{B}$). It took roughly 3 minutes to obtain this recommendation on a parallel computing server with 72 cores. While each posterior takes less than one second to approximate, we must approximate many posteriors to reliably characterize the uncertainty regarding which hypotheses are false. Our method could be expedited by considering multivariate analytical approximations to the posterior of $\boldsymbol{\theta}$. It would have taken over 20 minutes using the same computing resources to explore the sample size space using binary search instead of Algorithm 2; this discrepancy in runtime scales logarithmically with the recommended sample size.
@@
To confirm the strong performance of our method, we estimated the joint sampling distribution of posterior probabilities for $n_{A}=13328$ under the model $\Psi$ with $m=9.9\times 10^{4}$ simulation repetitions. Using a decision threshold of $\gamma=0.9411$, the FDR and average power in (1) were estimated as 0.0497 and 0.7992. This good performance occurs under the assumption that $\boldsymbol{\eta}\sim\Psi$. If $\Psi$ is misspecified such that the effect sizes are generally overstated, the true FDR and average power are likely to exceed $q$ and not exceed $1-\beta$, respectively. In contrast, a design informed by a model $\Psi$ that generally understates the effect sizes is likely to have an FDR and average power that outperform their respective criteria. We therefore advise that practitioners define $\Psi$ with respect to the minimum effect sizes they would like to detect.
@@
To demonstrate the flexibility of our method with different models and the value of considering the dependence structure at the design stage, we implemented Algorithm 2 with a modified version of $\Psi$ where the 5 five binary outcomes are generated independently according to the marginal probabilities from a given submodel $\Psi_{\boldsymbol{\Lambda}}$. The individual metrics in $\boldsymbol{\theta}$ were also analyzed using posteriors based on independent binomial models for each outcome and diffuse BETA(1,1) priors for each marginal probability. When ignoring the dependence structure between the 5 metrics, Algorithm 2 returned a recommended $(n_{A},\gamma)$ combination of $(14427,0.9500)$ under the constraint that $\boldsymbol{\gamma}=\gamma\times\mathbf{1}_{K}$. Thus, when accounting for the positive dependence between these binary outcomes, we can use a less strict decision threshold, resulting in a materially smaller sample size recommendation. If the outcomes exhibit negative dependence, we generally need a larger decision threshold – and sample size – to satisfy criteria for both the FDR and average power.
@@
Table 1 also demonstrates the value in considering more sophisticated optimization schemes to select $\boldsymbol{\gamma}$. For illustration, we also selected $\boldsymbol{\gamma}$ as the decision thresholds that ensure $\widehat{\text{FDR}}\leq q$ while maximizing average power subject to the constraint $\lvert\gamma_{k}-\gamma_{h}\rvert\leq 0.05$ for thresholds $k\neq h$. We detail our scheme to obtain the optimal $\boldsymbol{\gamma}_{n}$ for a given sample size $n$ in Appendix C.2. By optimizing the decision thresholds using more flexible constraints, we reduced the recommended sample size to $n_{A}=11106$. It took about 5 minutes on a parallel computing server to obtain this optimal design; 2 of these minutes were spent using a single core to implement our optimization scheme to find $\boldsymbol{\gamma}_{n}$ at each $n$ explored. While many optimization schemes not considered in this example could also be used to select $\boldsymbol{\gamma}$, the computational burden of the optimization scheme should be considered prior to implementing Algorithm 2.
Lastly, we compare our sample size recommendations to those prompted by a Bonferroni correction used to control the FWER of A/B tests in frequentist settings. To bound the FWER for 5 hypotheses by 0.05, the Bonferroni correction recommends using a significance level of 0.01 for each test. This significance level is similar to a decision threshold of $\gamma=0.99$ in our context. A sample size of $n_{A}=27956$ is recommended to ensure the average power across these 5 hypothesis tests is at least 80% while bounding the FWER by 0.05. We emphasize that control of the FDR and FWER are not interchangeable. Yet by comparing this sample size recommendation to those in Table 1, we illustrate that – even with as few as $K=5$ metrics – practitioners can reduce the required sample size for their A/B tests if willing to use designs that bound the FDR instead of the FWER. The benefit of our approach will be even more pronounced in cases where $K$ is very large. While it is rare for $K$ to be very large in online controlled experiments, our method could also be applied in contexts that typically involve a much larger number of comparisons.

### v5

@@ 3 A Proxy to the Joint Sampling Distribution
4 Economical Assessment of the FDR and Power
We generalize the results from Theorem 1 to develop the procedure in Algorithm 2. This procedure allows practitioners to efficiently explore the ($n,\boldsymbol{\gamma}$)-space to find the $(n,\boldsymbol{\gamma})$ combination under a given optimization scheme that minimizes the sample size while satisfying criteria for the Bayesian FDR and average power. Algorithm 2 is economical because we estimate the true joint sampling distribution of posterior probabilities via simulation at only two sample sizes: $n_{0}$ and $n_{1}$. The initial sample size $n_{0}$ is an input for Algorithm 2, and it can be selected based on the anticipated budget or timeline for the A/B test. We note that Algorithm 2 details a general application of our methodology, and we later describe potential modifications.
@@ 4 Economical Assessment of the FDR and Power
1:procedure Optimize(f()f($\cdot$), $g(\cdot)$, $\boldsymbol{\delta}_{L}$, $\boldsymbol{\delta}_{U}$, $p(\boldsymbol{\eta})$, $c$, $q$, $\beta$, $m$, $\Psi$, $n_{0}$)
@@
6:   end for
7:  end for
8:  Find $n_{1}$, the smallest $n$ such that Algorithm 1 with the expits of $\{\{\hat{l}^{{}_{(n)}}_{r,k}\}_{k=1}^{K}\}_{r=1}^{m}$ yields average power    $\geq 1-\beta$, where $\boldsymbol{\gamma}_{n}$ are the optimal thresholds such that $\widehat{\text{FDR}}$ $\leq q$
9:  Compute $\{\{\widehat{Pr}(H_{1,k}~|~\boldsymbol{y}^{{}_{(n_{1})}}_{r})\}_{k=1}^{K}\}_{r=1}^{m}$ via Algorithm 1 and their logits $\{\{\hat{l}^{{}_{(n_{1})}}_{r,k}\}_{k=1}^{K}\}_{r=1}^{m}$
10:  for $d$ in 11:$m$ do
11:   for $k$ in 11:$K$ do
12:     Let $\hat{l}^{{}_{(n_{0})}}_{d,k}$ and $\hat{l}^{{}_{(n_{1})}}_{d,k}$ be the $d^{\text{th}}$ order statistics of $\{\hat{l}^{{}_{(n_{0})}}_{r,k}\}_{r=1}^{m}$ and $\{\hat{l}^{{}_{(n_{1})}}_{r,k}\}_{r=1}^{m}$
13:     Let $r$ be the index of the sample $\boldsymbol{y}^{{}_{(n_{1})}}_{r}$ corresponding to $\hat{l}^{{}_{(n_{1})}}_{d,k}$
14:     Now use the line $\hat{L}^{{}_{(n)}}_{r,k}$ passing through $(n_{0},\hat{l}^{{}_{(n_{0})}}_{d,k})$ and $(n_{1},\hat{l}^{{}_{(n_{1})}}_{d,k})$ to get $\hat{l}^{{}_{(n)}}_{r,k}$ for other $n$
15:   end for
16:  end for
17:  Find $n_{2}$, the smallest $n$ such that Algorithm 1 with the expits of $\{\{\hat{l}^{{}_{(n)}}_{r,k}\}_{k=1}^{K}\}_{r=1}^{m}$ yields average power    $\geq 1-\beta$, where $\boldsymbol{\gamma}_{n}$ are the optimal thresholds such that $\widehat{\text{FDR}}$ $\leq q$
18:  return $n_{2}$ as recommended $n$ and $\boldsymbol{\gamma}_{n_{2}}$ as $\boldsymbol{\gamma}$
19:end procedure
@@
We choose the next sample size $n_{1}$ at which to estimate posterior probabilities in Line 8 of Algorithm 2. Recall that the optimal decision thresholds $\boldsymbol{\gamma}$ are linked with $n$ as described in Section 2. To account for this phenomenon, we run an optimization scheme to find the optimal $\boldsymbol{\gamma}$ for which $\widehat{\text{FDR}}$, the estimate of the Bayesian FDR defined in (4), is at most $q$. This optimization scheme can be as simple or as complex as the practitioner desires; we elaborate on several potential schemes in Section 5. Line 8 introduces a subscript $n$ for the decision thresholds to emphasize this dependence on the sample size. Given $\boldsymbol{\gamma}_{n}$ for a particular $n$, we can determine whether this sample size achieves an average power of at least $1-\beta$. The sample size $n_{1}$ is the smallest $n$ such that the average power is sufficiently large. We recommend finding this sample size using binary search. We emphasize that a modified version of Algorithm 1 based on the linear approximations from Line 5 of Algorithm 2 can be efficiently implemented for new sample sizes without computing additional posterior probabilities.
@@
In Lines 10 to 16 of Algorithm 2, we construct linear approximations to logits of posterior probabilities that are less reliant on large-sample results. These approximations are obtained separately for each metric $k$ by fitting a line through the same order statistic of the marginal sampling distribution at the sample sizes $n_{0}$ and $n_{1}$. Because Theorem 1 ensures that the limiting slopes of $\text{logit}\left(p^{{}_{(n)}}_{r,k}\right)$ and $\text{logit}\left(p^{{}_{(n)}}_{r,k}~|~\{p^{{}_{(n)}}_{r,h}\}_{h=1}^{k-1}\right)$ from the proxy sampling distribution are the same, we use the marginal sampling distributions to estimate slopes for the conditional logits. To maintain the proper level of dependence in the joint sampling distribution of posterior probabilities, we group the lines $\hat{L}^{{}_{(n)}}_{r,k}$ from Line 14 across metrics based on the sample $\boldsymbol{y}^{{}_{(n_{1})}}_{r}$ that defines the linear approximations. Given the linear trend in the proxy sampling distribution quantiles discussed in Section 3, it is reasonable to construct these linear approximations based on order statistics of estimates of the true sampling distributions when the $\boldsymbol{\theta}_{r}$ value is similar for all $\boldsymbol{\eta}_{r}\sim\Psi$.
@@
However, we recommend defining $\Psi$ as a mixture of various submodels, each of which gives rise to a different $\boldsymbol{\theta}_{r}$ value, to accommodate uncertainty about which hypotheses are false. This submodel approach is demonstrated in Section 5. In that case, the process in Lines 10 to 16 can be modified. We instead split the logits of the posterior probabilities for each $n$ value into subgroups based on their submodels in $\Psi$ before constructing the linear approximations. We use this set of linear approximations to obtain the optimal sample size $n_{2}$ and its associated decision thresholds in Line 17. Corollary 1 details how the Bayesian FDR and average power for the design defined using the $(n,\boldsymbol{\gamma})$ combination returned by Algorithm 2 respectively converge to $q$ and $1-\beta$ as the recommended value of $n\rightarrow\infty$.
Corollary 1.
Let the conditions for Theorem 1 be satisfied. Let the Bayesian FDR and average power be defined as in (1) and (3), respectively. Consider a design with the $(n,\boldsymbol{\gamma})$ combination recommended by Algorithm 2. As the recommended $n\rightarrow\infty$, this design is such that
$\left\lvert\mathbb{E}_{\boldsymbol{\eta}\sim\Psi}\left\{\mathbb{E}\left[\frac{V}{V+S}\right]\right\}-q\right\rvert\xrightarrow{P}0~~\text{and}~~\left\lvert\mathbb{E}_{\boldsymbol{\eta}\sim\Psi}\left\{\mathbb{E}\left[\frac{S}{T}\right]\right\}-(1-\beta)\right\rvert\xrightarrow{P}0.$
We prove Corollary 1 in Appendix C of the online supplement. While Algorithm 2 leverages asymptotic results, its linear approximations are constructed by simulating sampling distributions of posterior probabilities at finite sample sizes $n$. With finite sample sizes, Algorithm 2 thus recommends $(n,\boldsymbol{\gamma})$ combinations such that the Bayesian FDR approximates $q$ and average power approximates $1-\beta$. In Section 5, we consider the performance of this algorithm for an example based on a real A/B test involving 5 metrics.
@@ 5 Optimizely Example
We choose each set of multinomial probabilities for both groups using linear programming (Guenin et al., 2014). Linear programming allows us to construct multinomial models that satisfy the following objectives. First, the probabilities $\sum_{v=1}^{13}\eta_{j,v}$ must sum to 1 in both groups. Second, we must attain specified values for the marginal probabilities $\{\pi_{A,k}\}_{k=1}^{5}$ in group A. Third, we need the marginal probabilities $\{\pi_{B,k}\}_{k=1}^{5}$ in group B to achieve targets for lift based on anticipated values for $\{\theta_{k}\}_{k=1}^{5}$. These anticipated values determine which hypotheses in $\{H_{1,k}\}_{k=1}^{5}$ are true along with the effect sizes corresponding to the true hypotheses. We elaborate on this method to select a set of multinomial models in Appendix D.1 of the supplement.
@@
For a general A/B test, there are $K\choose k$ ways to classify $k$ of the $K$ hypotheses as false. In this Optimizely experiment, there are $\sum_{k=1}^{4}{5\choose k}=30$ unique classifications of $\{H_{1,k}\}_{k=1}^{5}$ such that some but not all of the hypotheses are false. We choose to define $\Psi$ by specifying 30 sets of multinomial models in Appendix D.1. We index these submodels by introducing a subscript to the notation $\Psi_{\boldsymbol{\Lambda}}$, where $\varnothing\subsetneq\boldsymbol{\Lambda}\subsetneq\{1,2,3,4,5\}$ is the set of hypotheses that are false. For the submodel $\Psi_{\boldsymbol{\Lambda}}$, the multinomial models for each group ensure that $\theta_{k}=0$ for $k\in\boldsymbol{\Lambda}$ and $\theta_{k}=10\%$ otherwise. It follows that $H_{1,k}$ is false in this submodel if and only if $k\in\boldsymbol{\Lambda}$. For illustration, we define the model $\Psi$ input into Algorithm 2 as an equally weighted mixture of these 30 submodels.
[…]
@@
Algorithm 2 with these inputs returned a recommended $(n_{A},\gamma)$ combination of $(13328,$ $0.9411)$. All sample size recommendations in this section pertain to the sample size per group (i.e., $n_{A}=n_{B}$). It took roughly 3 minutes to obtain this recommendation on a parallel computing server with 72 cores. While each posterior takes less than one second to approximate, we must approximate many posteriors to reliably characterize the uncertainty regarding which hypotheses are false. Our method could be expedited by considering multivariate analytical approximations to the posterior of $\boldsymbol{\theta}$. It would have taken over 20 minutes using the same computing resources to explore the sample size space using binary search instead of Algorithm 2; this discrepancy in runtime scales logarithmically with the recommended sample size.
@@
To confirm the strong performance of our method, we estimated the joint sampling distribution of posterior probabilities for $n_{A}=13328$ under the model $\Psi$ with $m=9.9\times 10^{4}$ simulation repetitions. Using a decision threshold of $\gamma=0.9411$, the Bayesian FDR and average power in (4) were estimated as 0.0497 and 0.7992. This good performance occurs under the assumption that $\boldsymbol{\eta}\sim\Psi$. If $\Psi$ is misspecified such that the effect sizes are generally overstated, the true Bayesian FDR and average power are likely to exceed $q$ and not exceed $1-\beta$, respectively. In contrast, a design informed by a model $\Psi$ that generally understates the effect sizes is likely to have a Bayesian FDR and average power that outperform their respective criteria. We therefore advise that practitioners define $\Psi$ with respect to the minimum effect sizes they would like to detect.
@@
To demonstrate the flexibility of our method with different models and the value of considering the dependence structure at the design stage, we implemented Algorithm 2 with a modified version of $\Psi$ where the 5 five binary outcomes are generated independently according to the marginal probabilities from a given submodel $\Psi_{\boldsymbol{\Lambda}}$. The individual metrics in $\boldsymbol{\theta}$ were also analyzed using posteriors based on independent binomial models for each outcome and diffuse BETA(1,1) priors for each marginal probability. When ignoring the dependence structure between the 5 metrics, Algorithm 2 returned a recommended $(n_{A},\gamma)$ combination of $(14427,0.9500)$ under the constraint that $\boldsymbol{\gamma}=\gamma\times\mathbf{1}_{K}$. Thus, when accounting for the positive dependence between these binary outcomes, we can use a less strict decision threshold, resulting in a materially smaller sample size recommendation. If the outcomes exhibit negative dependence, we generally need a larger decision threshold – and sample size – to satisfy criteria for both the Bayesian FDR and average power.
@@
Table 1 also demonstrates the value in considering more sophisticated optimization schemes to select $\boldsymbol{\gamma}$. For illustration, we also selected $\boldsymbol{\gamma}$ as the decision thresholds that ensure $\widehat{\text{FDR}}\leq q$ while maximizing average power subject to the constraint $\lvert\gamma_{k}-\gamma_{h}\rvert\leq 0.05$ for thresholds $k\neq h$. We detail our scheme to obtain the optimal $\boldsymbol{\gamma}_{n}$ for a given sample size $n$ in Appendix D.2. By optimizing the decision thresholds using more flexible constraints, we reduced the recommended sample size to $n_{A}=11106$. It took about 5 minutes on a parallel computing server to obtain this optimal design; 2 of these minutes were spent using a single core to implement our optimization scheme to find $\boldsymbol{\gamma}_{n}$ at each $n$ explored. While many optimization schemes not considered in this example could also be used to select $\boldsymbol{\gamma}$, the computational burden of the optimization scheme should be considered prior to implementing Algorithm 2.
Lastly, we compare our sample size recommendations to those prompted by a Bonferroni correction used to control the FWER of A/B tests in frequentist settings. To bound the FWER for 5 hypotheses by 0.05, the Bonferroni correction recommends using a significance level of 0.01 for each test. This significance level is similar to a decision threshold of $\gamma=0.99$ in our context. A sample size of $n_{A}=27956$ is recommended to ensure the average power across these 5 hypothesis tests is at least 80% while bounding the FWER by 0.05. We emphasize that control of the (Bayesian) FDR and FWER are not interchangeable. Yet by comparing this sample size recommendation to those in Table 1, we illustrate that – even with as few as $K=5$ metrics – practitioners can reduce the required sample size for their A/B tests if willing to use designs that bound the Bayesian FDR instead of the FWER. The benefit of controlling the Bayesian FDR will be even more pronounced in cases where $K$ is very large. While it is rare for $K$ to be very large in online controlled experiments, our method could also be applied in contexts that typically involve a much larger number of comparisons.